# Geometric structure of shallow neural networks and constructive ${\mathcal{L}}^{2}$ cost minimization

### v3

@@ 1. Introduction
Applications of neural networks have rapidly become indispensable in a vast range of research disciplines, and their technological impact in recent years has been enormous. However, despite those successes, the fundamental conceptual reasons underlying their functioning are currently insufficiently understood, and remain the subject of intense investigation, see e.g., [LBH15, Aro18, GK23]. The common approach to the minimization of the cost function is based on the gradient flow generated by it, and a variety of powerful algorithms have been developed for this purpose that can be highly successful in suitable application scenarios; nevertheless, even if a global or sufficiently good local minimum is obtained, the properties of the minimizing parameters (weights, biases) in general remain quite mysterious, see e.g., [HR18, NRS21].
@@
Finally, a comment on the assumptions: We study shallow neural networks with dimensions of input, hidden and output layer $(M,M,Q)$, respectively, where $M\geq Q$ and at one point specialize to $M=Q$. While this is restrictive in the context of general neural networks, this can be useful as the study of the last layers of a deep network. Moreover, the current paper is the first in a series of works in which we investigate the geometric structure of cost minimizers in neural networks, including deep networks with more general architectures.
@@ 1.1. Related work
Despite the power of deep networks, theory of shallow networks is still actively studied. Examples of works on shallow networks include [MVEZ20, ABK+25].
For the ReLU activation function, [Kaw16] and [ZL18] studied minimizers by simplifying to linear networks, with no bias vectors. [Coo21] makes use of the bias vectors to construct a shallow ReLU network which interpolates any data; however, the resulting network is highly overparametrized.
For linear networks, the classic work [BH89] studied critical points of the square loss for shallow linear networks. This was extended in [Kaw16] to deep linear networks, and further studied in works such as [ZL18, TKB19, AMG24].
@@
We present the explicit construction of local and global ${\mathcal{L}}^{2}$ cost minimizers in underparametrized deep (that is, multilayer) networks in a separate work [CE23], which invokes results developed in the paper at hand, combined with inspiration drawn from rigorous renormalization group analysis in quantum field theory, as for instance in [BCFS03].
@@ 1.2. Outline of paper
This paper is organized as follows. In Section 2, we give a detailed description of the mathematical model and introduce notation. In Section 3, we present the main results, namely Theorem 3.1, which we prove in Section 4, Theorem 3.2, which we prove in Section 5, Theorem 3.3, which we prove in Section 6, and Theorem 3.5.
@@ 2. Definition of the mathematical model
$B_{1}$ ::=:= $b_{1}u_{N}^{T}=[b_{1}\cdots b_{1}]\in{\mathbb{R}}^{M\times N}$
$B_{2}$ ::=:= $b_{2}u_{N}^{T}=[b_{2}\cdots b_{2}]\in{\mathbb{R}}^{Q\times N}$ (2.2)
@@
[…]
@@ 3.1. Upper bound on minimum of cost function for $M>Q$
On the other hand, the $M-Q$ rows of $RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*}$ are eliminated by way of
@@
Let $Q\leq M\leq QM<N$. Assume that $R\in O(M)$ diagonalizes $P,P^{\perp}$, and let $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$.
@@ 3.2. Exact degenerate local minimum in the case $M=Q$
Assume $M=Q<MQ<N$. Then, if ${\rm rank}(\overline{X_{0}})=Q$, let
@@ 3.3. Geometric interpretation
Assume $Q\leq M\leq QM<N$. Let $W_{i}^{*},b_{i}^{*}$, $i=1,2$, denote the weights and biases determined in Theorem 3.1. Let
@@
on the $Q$-dimensional linear subspace ${\rm range}(P)\subset{\mathbb{R}}^{M}$, where ||$\cdot$|| ⋅ | denotes the Euclidean norm on ${\mathbb{R}}^{Q}$. Then,
@@ 3.4. Dependence on truncation
$\Delta_{2}^{rel,tr}$ ::=:= $\Delta_{1}^{rel,tr}{\mathcal{N}}^{-1}(\Delta_{1}^{rel,tr})^{T}\;\;,\;{\rm with}\;$
$\Delta_{1}^{rel,tr}$ ::=:= $(\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}})^{-1}\Delta(\tau_{W_{1},b_{1}}(x_{0,j,i}))\,,$ (3.48)
@@
where $P_{R}$ has $Q$ diagonal entries of 111, and all other components are 0, while $P^{\perp}_{R}$ has $M-Q$ diagonal entries of 111, and all other components are 0. This does not characterize $R$ uniquely, as we may compose $R$ with an arbitrary $R^{\prime}\in O(M)$ that leaves the ranges of $P,P^{\perp}$ invariant. Diagonality of $P_{R},P_{R}^{\perp}$ implies that the activation function, which acts component-wise as the ramp function (2.5), satisfies
@@
is a matrix all of whose components are 0$\leq$ 0≤ 0. Therefore, $\sigma$ maps it to 0.
@@
0 0 $=$ $\partial_{b_{2}}\|W_{2}RP\Delta X_{0}+\beta_{2}u_{N}^{T}\|_{{\mathcal{L}}^{2}}^{2}$ (4.32)
@@
${\mathcal{R}}$ ::=:= $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$ (5.31)
@@
0 0 $=$ $\partial_{b_{2}}\|Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ (5.34)
@@
$\Delta_{1}^{rel}$ ::=:= $(\overline{X_{0}^{red}})^{-1}\Delta X_{0}$
$\Delta_{2}^{rel}$ ::=:= $(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](\overline{X_{0}^{red}})^{-T}$ (5.37)
@@
$(I)$ ::=:= ${\rm Tr}\Big{(}Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}(1+\Delta_{2}^{rel})^{-T}Y^{T}\Big{)}$ (5.41)
@@
$(II)$ ::=:= $-2{\rm Tr}\Big{(}Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}\Delta_{2}^{rel}Y^{T}\Big{)}$ (5.42)
@@
$(III)$ ::=:= ${\rm Tr}\Big{(}Y\Delta_{2}^{rel}\big{(}1+\Delta_{2}^{rel}\big{)}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}$ (5.43)

### v4

# arXiv:2309.10370v4[cs.LG] 26 Feb 2026

@@ 1. Introduction
Applications of neural networks have rapidly become indispensable in a vast range of research disciplines, and their technological impact in recent years has been enormous. However, despite those successes, the fundamental conceptual reasons underlying their functioning are currently insufficiently understood, and remain the subject of intense investigation. The common approach to the minimization of the cost function is based on the gradient flow generated by it, and a variety of powerful algorithms have been developed for this purpose that can be highly successful in suitable application scenarios, see e.g. [LBH15, Aro18, GK23]; nevertheless, even if a global or sufficiently good local minimum is obtained, the properties of the minimizing parameters (weights, biases) in general remain quite mysterious, see e.g., [HR18, NRS21].
@@
In Section 4, we compare our theoretical bounds with numerical results based on randomly initialized neural networks trained with gradient descent on synthetic data. We show that the bound improves with the increase of clustering of the data, whereby $\delta_{P}$ decreases; this matches our prediction.
Finally, we comment on the assumptions: We study shallow neural networks with dimensions of input, hidden and output layer $(M,M,Q)$, respectively, where $M\geq Q$ and at one point specialize to $M=Q$. While this is restrictive in the context of general neural networks, this can be useful as the study of the last layers of a deep network.
Moreover, we place no restriction on the number of training data points $N$ compared with the number of parameters, and so the results here hold for both underparametrized and overparametrized networks. In the strongly overparametrized regime $M>N$, every data point is its own cluster, so that the cluster variances are identically zero. Therefore, our construction yields zero training loss. This situation and its extension to deep networks is studied in detail in [BCC+25].
@@ 1.1. Related work
Early work on the loss landscape of neural networks already focused on shallow architectures. For linear networks, [BH89] analyzed the square loss for one-hidden-layer linear networks and showed that, under mild assumptions, every local minimum is global. [Kaw16] extended this to deep linear networks, proving the absence of bad local minima and characterizing saddle points, and subsequent works refined the description of critical points and their geometry, including explicit parametrizations and second-order analyses [ZL18, TKB20, AMG24]. For shallow ReLU networks, the loss landscape can be partly reduced to a linear problem when biases are ignored, and this reduction was exploited in [Kaw16, ZL18]. [Coo21] instead made crucial use of bias vectors to construct global minima that lie outside a fixed linear region for overparameterized networks, thereby showing that genuinely nonlinear structure in the loss landscape is already present in the shallow case.
Despite the dominance of very deep architectures in applications, shallow networks remain a central object in contemporary theory. From the optimization and generalization side, mean-field analyses for two-layer networks describe gradient-based training via limiting partial differential equations and yield convergence and generalization guarantees, e.g. [MMN18, MMM19]. In parallel, a substantial literature studies the implicit bias of gradient descent and gradient flow, both in regression and classification, e.g. [SHN+18, GLSS18, NCS19, LL20, JM23].
Shallow models also appear implicitly when one “freezes” all but the last few layers of a deep network. This perspective underlies work on linear classifier probes, where linear predictors are trained on top of intermediate representations to study how separability emerges across layers [AB17], and it is also closely connected to the unconstrained features model (UFM) and the phenomenon of neural collapse. In the UFM, one replaces the learned features by free variables and optimizes the last-layer problem jointly over logits and features, reducing a deep model to a structured shallow one [ZDZ+21, EW22, MPP22, TB22, HL24]. In all these settings, shallow architectures serve both as analytically tractable proxies and as faithful models of the final stages of deep networks.
Finally, shallow networks are also a minimal setting in which one can contrast lazy-training (kernel-like) and feature-learning regimes. In the infinite-width limit with appropriate scaling, gradient descent on two-layer ReLU networks evolves essentially as a linear model around initialization and leaves the representation almost fixed, the so-called lazy-training regime [JGH18, LXS+19, COB19]. By changing width, initialization scale, or training time one exits this regime and enters a genuinely nonlinear feature-learning regime where the representation itself evolves during training [LXMZ21]. Together with mean-field analyses that capture representation dynamics, this shows that shallow ReLU networks alone realize both extremes and the interpolation between them [MU25], while remaining simple enough that their loss landscape and implicit bias can be analyzed in detail.
1.2. Related work by the authors
We present the explicit construction of local and global ${\mathcal{L}}^{2}$ cost minimizers in deep neural networks for clustered data in a separate work [CE23], which invokes results developed in the paper at hand, combined with inspiration drawn from rigorous renormalization group analysis in quantum field theory, as for instance in [BCFS03]. In [CME25], we extend this construction to sequentially linearly separable data sets which are significantly less structured than in [CE23].
In Section 3.4, we address the structure of critical sets for shallow networks. More generally, the comparison of constructive training error bounds with those obtained from gradient descent is an interesting theoretical problem. A detailed analysis of the dynamical systems aspects of gradient descent and its geometric interpretation for both shallow and deep networks is given in [Che23, CME24].
@@ 1.3. Outline of paper
This paper is organized as follows. In Section 2, we give a detailed description of the mathematical model and introduce notation. In Section 3, we present the main results, namely Theorem 3.1, which we prove in Section 5, Theorem 3.2, which we prove in Section 6, Theorem 3.3, which we prove in Section 7, and Theorem 3.5. In Section 4, we present some numerical experiments.
@@ 2. Definition of the mathematical model
$B_{1}$ $:=$ $b_{1}u_{N}^{T}=[b_{1}\cdots b_{1}]\in{\mathbb{R}}^{M\times N}$
$B_{2}$ $:=$ $b_{2}u_{N}^{T}=[b_{2}\cdots b_{2}]\in{\mathbb{R}}^{Q\times N}$ (2.2)
@@
and we note that ${\rm Pen}[\overline{X_{0}^{red}}]\overline{X_{0}^{red}}={\bf 1}_{Q\times Q}$. The projector property $P^{2}=P$ is thus easily verified, and orthogonality with respect to the Euclidean inner product $(\cdot,\cdot)$ on ${\mathbb{R}}^{M}$ holds due to $P^{T}=P$, whereby $(v,Pw)=(Pv,w)$ for any $v,w\in{\mathbb{R}}^{M}$. In particular, we have
@@ 3.1. Upper bound on minimum of cost function for $M>Q$
On the other hand, the MQM-Q rows of $RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*}$ are eliminated by way of
@@
Let $Q\leq M\leq QM$. Assume that $R\in O(M)$ diagonalizes $P,P^{\perp}$, and let $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$.
@@ 3.2. Exact degenerate local minimum in the case $M=Q$
Assume $M=Q<MQ$. Then, if ${\rm rank}(\overline{X_{0}})=Q$, let
@@ 3.3. Geometric interpretation
Assume $Q\leq M\leq QM$. Let $W_{i}^{*},b_{i}^{*}$, $i=1,2$, denote the weights and biases determined in Theorem 3.1. Let
@@
on the $Q$-dimensional linear subspace ${\rm range}(P)\subset{\mathbb{R}}^{M}$, where $|\cdot|$ denotes the Euclidean norm on ${\mathbb{R}}^{Q}$. Then,
@@ 3.4. Dependence on truncation
$\Delta_{2}^{rel,tr}$ $:=$ $\Delta_{1}^{rel,tr}{\mathcal{N}}^{-1}(\Delta_{1}^{rel,tr})^{T}\;\;,\;{\rm with}\;$
$\Delta_{1}^{rel,tr}$ $:=$ $(\overline{(\tau_{W_{1},b_{1}}(X_{0}))^{red}})^{-1}\Delta(\tau_{W_{1},b_{1}}(x_{0,j,i}))\,$, (3.48)
@@
4. Experiments
To assess the validity of our theoretical results in practical training scenarios, we generated synthetic data using a Gaussian mixture model in $\mathbb{R}^{M}$ with $Q$ classes, each having randomly distributed means and varying standard deviations. For each dataset, constructed with a fixed standard deviation and $Q$ classes, we initialized and trained a ReLU neural network with architecture $(M,M,Q)$, as defined in (2.6)-(2.7), to classify the data according to the cost function (2.27). Each configuration was trained 100 times. Figures 1 and 2 display the average final cost across runs, the corresponding average initial cost, and the bound (3.15) from Theorem 3.1, which we repeat here:
$\text{Bound}=\frac{1}{\sqrt{N}}\left\lVert{\left(\overline{X_{0}^{red}}\right)^{+}\Delta X_{0}}\right\rVert_{\mathcal{L}^{2}}.$ (4.1)
Note that we taking the label vectors to be the canonical basis vectors in $\mathbb{R}^{Q}$, so that $Y$ is the identity.
Figure 1. (Average of) initial ($C_{init}$) and final ($C_{final}$) cost of randomly initialized neural network(s) trained to classify Gaussian mixture data with $Q$ classes, for different (fixed in each plot) cluster standard deviations, plotted against the bound (4.1) computed for each data set.
Figure 1 shows that as the variance of the clusters in the Gaussian mixture decreases, the theoretical bound increasingly aligns with the average final cost achieved during training. In particular, in Figure 2, corresponding to the smallest variance, the bound closely tracks the final training loss and, for some runs, is even smaller than the attained cost.
Figure 2. (Average of) initial ($C_{init}$) and final ($C_{final}$) cost of randomly initialized neural network(s) trained to classify Gaussian mixture data with $Q$ classes, plotted against the bound (4.1) computed for each data set.
We summarize the experiment parameters, as well as the hyperparameters used in the training of the neural networks via stochastic gradient descent, in Table 1. Code is available at https://github.com/patriciaewald/bounds.
Hyperparameter Notation Value
Input dimension $M$ $10$
Samples per class $N_{j}$ $100$
Number of classes $Q$ $\{10,\cdots,2\}$
Cluster standard deviation std $\{0.5,\cdots,0.1\}$
Random initializations $100$
Maximum number of epochs $1000$
Learning rate $0.05$
Momentum $0.9$
Plateau tolerance $10^{-4}$
Plateau patience (epochs) $30$
Table 1. Hyperparameters used in the Gaussian mixture and network training experiments.
@@ 4. Experiments
where $P_{R}$ has $Q$ diagonal entries of $1$, and all other components are 0, while $P^{\perp}_{R}$ has MQM-Q diagonal entries of $1$, and all other components are 0. This does not characterize $R$ uniquely, as we may compose $R$ with an arbitrary $R^{\prime}\in O(M)$ that leaves the ranges of $P,P^{\perp}$ invariant. Diagonality of $P_{R},P_{R}^{\perp}$ implies that the activation function, which acts component-wise as the ramp function (2.5), satisfies
@@
is a matrix all of whose components are $\leq 0$. Therefore, $\sigma$ maps it to 0.
@@
$0$ $=$ $\partial_{b_{2}}\|W_{2}RP\Delta X_{0}+\beta_{2}u_{N}^{T}\|_{{\mathcal{L}}^{2}}^{2}$ (5.32)
@@
${\mathcal{R}}$ $:=$ $Y_{b_{2}}(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](X_{0}{\mathcal{N}}^{-1}X_{0}^{T})^{-1}X_{0}$ (6.31)
@@
$0$ $=$ $\partial_{b_{2}}\|Y^{ext}_{b_{2}}{\mathcal{P}}^{\perp}\|_{{\mathcal{L}}^{2}_{{\mathcal{N}}}}^{2}$ (6.34)
@@
$\Delta_{1}^{rel}$ $:=$ $(\overline{X_{0}^{red}})^{-1}\Delta X_{0}$
$\Delta_{2}^{rel}$ $:=$ $(\overline{X_{0}^{red}})^{-1}D_{2}[X_{0}](\overline{X_{0}^{red}})^{-T}$ (6.37)
@@
$(I)$ $:=$ ${\rm Tr}\Big(Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}(\Delta_{1}^{rel})^{T}(1+\Delta_{2}^{rel})^{-T}Y^{T}\Big)$ (6.41)
@@
$(II)$ $:=$ $-2{\rm Tr}\Big(Y(1+\Delta_{2}^{rel})^{-1}\Delta_{1}^{rel}{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}\big(1+\Delta_{2}^{rel}\big)^{-1}\Delta_{2}^{rel}Y^{T}\Big)$ (6.42)
@@
$(III)$ $:=$ ${\rm Tr}\Big(Y\Delta_{2}^{rel}\big(1+\Delta_{2}^{rel}\big)^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]{\mathcal{N}}^{-1}[e_{1}u_{N_{1}}^{T}\cdots e_{Q}u_{N_{Q}}^{T}]^{T}$ (6.43)